Ecob Corporation uses the absorption costing approach to cost-plus pricing as described in the text to set prices for its products. Based on budgeted sales of 32,000 units next year, the unit product cost of a particular product is $61.80. The companys selling and administrative expenses for this product are budgeted to be $387600 in total for the year. The company has invested $400,000 in this product and expects a return on investment of 11%. The markup on absorption cost for this product would be closest to: (Do not round intermediate calculations) 05903 Multiple Choice 174% 21 8% 11.0%
Answer #1
Required ROI = 400000*11%= $44000
Markup amount per unit = (387600+44000)/32000= $13.49
Markup percentage on absorption cost = 13.49/61.80= 21.8%
Option 2 is correct